OpenStudy (anonymous):

Stainless steel is a corrosion resistant alloy composed of nickel, chromium and iron. A lab technician was assigned to determine the mass of chromium in a 2.50 g sample of stainless steel. The sample was first dissolved in acid to convert the chromium metal in the sample into chromium (II) ions. The acidified chromium (II) ions were then titrated with 17.69 mL of 0.125 mol/L perchlorate solution according to the unbalanced equation: ClO4- +Cr2+ --> Cl- + Cr 3+ Calculate the mass of chromium found in the 2.50 g sample of stainless steel.

myininaya (myininaya):

Is there an equation to use?

myininaya (myininaya):

a mathematical one

OpenStudy (anonymous):

yes

myininaya (myininaya):

do we need to balance the chemical equation first?

myininaya (myininaya):

wait i don't see how

OpenStudy (anonymous):

i think so.. half reactions?

myininaya (myininaya):

there are no O on the other side

OpenStudy (anonymous):

oxidation and reduction half reactions (redox)

myininaya (myininaya):

was trying to find an example

myininaya (myininaya):

i think this is definitely beyond my chemistry understanding i'm sorry i cannot help

OpenStudy (anonymous):

it's okay... thanks for trying ! :)

OpenStudy (anonymous):

Do you know anyone who is good at chemistry?

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=balance+ClO4-+%2BCr2%2B+--%3E+Cl-+%2B+Cr+3%2B I don't know if this helps... But it says the molar mass in chromuim(II) is 51.9961 g/mol and chromuim(III) is 51.9961 g/mol

myininaya (myininaya):

I don't know if molar mass is different during regular mass or what the relationship is or how they got those numbers but maybe somehow you can use it to check any work you have done so far

myininaya (myininaya):

and for some reason it couldn't balance it

OpenStudy (anonymous):

I think you need to separate them into half reactions by using "table of selected Standard Electrode Potentials"

OpenStudy (anonymous):

than balancing the electrons and put them in a net equation...

myininaya (myininaya):

I have to read up on this. I will be back tomorrow.

myininaya (myininaya):

I sorta found an example but it is with balanced chemical equations

myininaya (myininaya):

http://wps.prenhall.com/wps/media/objects/3311/3390507/blb0406.html So I'm going to read this...

myininaya (myininaya):

\[17.69 mL \cdot \frac{1 L}{1000 mL} \cdot ? \frac{mol Cr }{L}\]

myininaya (myininaya):

I don't know what that question mark should be though

myininaya (myininaya):

we are not given the mol/L thing for Cr

OpenStudy (anonymous):

molar ratio for the question mark

OpenStudy (anonymous):

I find this link... http://kcourses.weebly.com/uploads/1/6/8/1/16813808/2.5-redox_stoichometry.pdf

OpenStudy (sweetburger):

i think this is a redox titration involving some gravimetrical analyses

OpenStudy (anonymous):

yes, it is. im not sure how to approach it.

OpenStudy (anonymous):

I know that you need to make half reactions...

OpenStudy (sweetburger):

alright first determine the moles of percholorate .01769L X .125 mole/1L

OpenStudy (anonymous):

it is 0.00221125mol